OpenStudy (anonymous):

How do you find the length of the focal chord for: y^2+6y-2x+13=0 ?

OpenStudy (mertsj):

put it into the form y^2=4px 4p is the length of the focal chord.

OpenStudy (anonymous):

I got 1, is that right? @Mertsj

OpenStudy (mertsj):

I got 2

OpenStudy (mertsj):

y^2+6y+13=2x (y+3)^2=2x-4 (y+3)^2=2(x-4) 4p=2